OpenStudy (anonymous):

f(x)=x^2n / (2n)! . what is f'(x) ??

OpenStudy (lalaly):

\[f'(x)=\frac{2nx^{2n-1}}{2n!}=\frac{nx^{2n-1}}{n!}\]

OpenStudy (anonymous):

here is the \[\sum_{n=0}^{\infty} f(x)\]

OpenStudy (anonymous):

weird the answer should be \[x ^{2n+1}/(2n+1)!\]